# Spatial Sign based Principal Component Analysis for High Dimensional Data

### v1

@@ 7.2.1 Some useful lemmas
$\hat{\boldsymbol{U}}_{i}-\boldsymbol{U}_{i}=$ $\frac{\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}}{\|\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}\|}-\frac{\boldsymbol{X}_{i}-\boldsymbol{\mu}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|}$
$=$ $\frac{\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}}{\|\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}\|}-\frac{\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|}+\frac{\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|}-\frac{\boldsymbol{X}_{i}-\boldsymbol{\mu}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|}$
$=$ $(\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}})(\hat{r}_{i}^{-1}-{r}_{i}^{-1})+(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})r_{i}^{-1}$
$=$ $\boldsymbol{U}_{i}(r_{i}\hat{r}_{i}^{-1}-1)+(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})(\hat{r}_{i}^{-1}-{r}_{i}^{-1})+(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})r_{i}^{-1}.$
@@
Because $|\hat{r}_{i}-r_{i}|\leq\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|$, so $|r_{i}\hat{r}_{i}^{-1}-1|\leq\frac{\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|}{r_{i}-\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|}$. Then,
$\max_{1\leq i\leq T}|r_{i}\hat{r}_{i}^{-1}-1|\leq\zeta_{1}\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|\max_{1\leq i\leq T}(\zeta_{1}r_{i}-\zeta_{1}\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|)^{-1}.$
By Lemma 7.4, we have $\zeta_{1}\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|=O_{p}(n^{-1/2})$. And by the sub-gaussian assumption of $\nu_{i}$, we have $\max_{1\leq i\leq n}\nu_{i}=O_{p}(\log^{1/2}n)$. So $\max_{1\leq i\leq T}|r_{i}\hat{r}_{i}^{-1}-1|=O_{p}(\sqrt{\frac{\log n}{n}})$. In addition $\|\tilde{\bf S}\|_{max}\leq\|\tilde{{\bf S}}-{\bf S}\|_{max}+\|{\bf S}\|_{max}$. So, with probability larger than $1-\alpha$,

### v2

@@ 6 Conclusion
$\hat{\boldsymbol{U}}_{i}-\boldsymbol{U}_{i}=$ $\frac{\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}}{\|\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}\|{{}_{2}}}-\frac{\boldsymbol{X}_{i}-\boldsymbol{\mu}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|{{}_{2}}}$
$=$ $\frac{\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}}{\|\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}\|{{}_{2}}}-\frac{\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|{{}_{2}}}+\frac{\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|{{}_{2}}}-\frac{\boldsymbol{X}_{i}-\boldsymbol{\mu}}{\|\boldsymbol{X}_{i}-\boldsymbol{\mu}\|{{}_{2}}}$
$=$ $(\boldsymbol{X}_{i}-\hat{\boldsymbol{\mu}})(\hat{r}_{i}^{-1}-{r}_{i}^{-1})-(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})r_{i}^{-1}$
$=$ $\boldsymbol{U}_{i}(r_{i}\hat{r}_{i}^{-1}-1)-(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})(\hat{r}_{i}^{-1}-{r}_{i}^{-1})-(\hat{\boldsymbol{\mu}}-\boldsymbol{\mu})r_{i}^{-1}.$
@@
Because $|\hat{r}_{i}-r_{i}|\leq\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|{{}_{2}}$, so $|r_{i}\hat{r}_{i}^{-1}-1|\leq\frac{\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|{{}_{2}}}{r_{i}-\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|{{}_{2}}}$. Then,
$\max_{1\leq i\leq T}|r_{i}\hat{r}_{i}^{-1}-1|\leq\zeta_{1}\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|{{}_{2}}\max_{1\leq i\leq T}(\zeta_{1}r_{i}-\zeta_{1}\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|{{}_{2}})^{-1}.$
By Lemma S2.4, we have $\zeta_{1}\|\hat{\boldsymbol{\mu}}-\boldsymbol{\mu}\|{{}_{2}}=O_{p}(n^{-1/2})$. And by the sub-gaussian assumption of $\nu_{i}$, we have $\max_{1\leq i\leq n}\nu_{i}=O_{p}(\log^{1/2}n)$. So $\max_{1\leq i\leq T}|r_{i}\hat{r}_{i}^{-1}-1|=O_{p}(\sqrt{\frac{\log n}{n}})$. In addition $\|\tilde{\bf S}\|_{max}\leq\|\tilde{{\bf S}}-{\bf S}\|_{max}+\|{\bf S}\|_{max}$. So, with probability larger than $1-\alpha$,